Analyze Dollar General Dollar General Corporation (DG) is a discount retailer with more than 12,000 stores. It offers a wide.................
Answer #1

a. Assets turnover = Total sales / average accounts receivable

Year 2 = 21987/(11258+11672)/2 = 1.92

Year 1 = 20369/(11209+11258)/2 = 1.81